A computer chess game and a human chess champion are evenly matched. They play fourteen games. Find probabilities for the following events. (Round all answers to four decimal places.)
(a) They each win seven games.
P =
(b) The human chess champion wins ten games.
P =
(c) The computer wins at least ten games.
P =
Answer #1

a) P ( X =    7   ) = C(14,7) * 0.5^7 * (1-0.5)^7 =            0.2095   (answer)

b) P ( X =    10   ) = C(14,10) * 0.5^10 * (1-0.5)^4 =            0.0611   (answer)

c) P(X≥10) = Σ14C x * 0.5^ x * (1-0.5)^(14-x) where x varies from 10 to 14 = 0.0898